# Hybrid plasmonic nanoantennas using corner reflectors for enhanced single-photon emission

### v1

@@ Abstract
The emission rate of atom-like photon sources can be significantly improved by coupling them to plasmonic resonant nanostructures. These arrangements function as nanoantennas, serving the dual purpose of enhancing light-matter interactions and decoupling the emitted photons. However, there is a contradiction between the requirements for these two tasks. A small resonator volume is necessary for maximizing interaction efficiency, while a large antenna mode volume is essential to achieve high emission directivity. In this work, we analyze a hybrid structure composed of a noble metal plasmonic resonant nanoparticle coupled to the atom-like emitter, designed to enhance the emission rate, alongside a corner reflector aimed at optimizing the angular distribution of the emitted photons. A comprehensive numerical study of silver and gold corner reflector nanoantennas, employing the finite-difference time-domain method, is presented. The results demonstrate that a well-designed corner reflector can significantly enhance photon emission directivity while also substantially boosting the emission rate.
@@ 1 Introduction
Efficient and reliable single-photon sources, capable of producing on-demand high-purity single-photon states, are key elements for the emerging field of photon-based quantum technologies [1, 2]. Atom-like emitters, based either on solid-state or molecular systems, are a promising alternative for this purpose [3, 4, 5, 6, 7]. In the past years, nanoscale elements have been introduced to enhance the emission rate of the photon sources and to tailor their spatial radiation properties. Such nanoantennas (NAs) reproduce at the nanoscale the action of traditional radiofrequency (RF) aerials. Various traditional antenna geometries have been considered for applications in the optical realm: dipole [8, 9], patch [10], bowtie [11] or the Yagui-Uda array [12, 13, 14].
@@
Antennas for light [12] serve the dual purpose of enhancing the emitter fluorescence and facilitating the out-coupling of the generated photons. The performance requisites for the two functions are antithetical: whereas a small mode volume is required for an enhanced emission rate, high antenna directivity is well known to be the result of a large electrical size, i.e, the extent of the structure measured using the wavelength of the electromagnetic field as unit reference [15, 16]. In [4], such antipodal requirements have been reconciled by photomodifying a nanopatch antenna assembled around a nanodiamond with a nitrogen-vacancy center to reduce the the cavity mode size while keeping a large size emitter to spatially shape the radiation produced. In this work, we tackle this issue by using a hybrid plasmonic structure built from the combination of a conventional NA acting as the feeder element of a corner reflector (CR) [17]. This strategy permits simultaneously optimize the emission rate at the feeder design and to use a CR to enhance the directivity of the photons produced.
@@
The use of CRs in antenna design has been prevalent since Kraus introduced them in 1940 [17], owing to their outstanding performance, simplicity, and ease of construction. This work demonstrates that these advantageous properties can also be applied to nanoscale photonics. The potential of CRs to enhance photon emission was highlighted in [18], where the radiation patterns generated by an electric dipole emitter with a perfect conductor CR were analyzed. Here, we present a thorough numerical study of the proposed NA, utilizing two alternative noble metals: silver and gold. The feeder is assumed to be constituted by an atom-like dipole emitter coupled to a spherical nanoparticle (NP), though the proposed CRs can also accommodate more complex feeder geometries, such as dimers or bowtie structures [8, 11], which may offer enhanced emission. The analysis is conducted using the finite-difference time domain method (FDTD)[19, 20], enabling the calculation of the radiative emission rate enhancement factor, losses within the metallic structures, and the far-field patterns of the emitted photons. This approach provides a comprehensive characterization of the NA.
@@
The geometry employed in the analyses is depicted in Fig. 1. A spherical NP with diameter $d$ is positioned at the origin. The atom-like emitter is represented by a point dipole source $\mathbf{p}$ located at a distance $s$ from the surface of the metal nanosphere that is placed at a distance $S$ from the corner apex along the bisector. The corner angle is $\psi$, while the metal films have a length $L$ and height $H$, as shown in the figure. The thickness of the metal films forming the reflector is $W$. The planes of the CR are assumed to be made of the same metal as used for the NP in the feeder system.
@@
A metallic plane typically reflects a substantial portion of an incident wave originating from a large distance, due to the significant mismatch between its optical properties and those of the surrounding dielectric. However, when a noble metal plane is placed near a dipolar source, it can significantly influence the emission rate [21], similar to the behavior observed with emission near NPs. Therefore, analyzing only the radiation pattern of emitted photons is insufficient for designing CR-based NAs. A comprehensive analysis, including the effects on radiative emission and loss rates, has been conducted. The impact of the thickness of the metal films on these properties has also been addressed.
@@ 2 Materials and Methods
The numerical calculations were conducted using ANSYS Lumerical FDTD Solutions, release 2024 R2.1, on a HP Z6 G4 Workstation with two Intel Xeon 6226R processors and 512GB of RAM. For the finite-difference time-domain (FDTD) simulations, a variable-size discretization was employed, with a finer grid in the dipole-nanoparticle region. The discretization parameters were adjusted based on a series of convergence analyses, setting the maximum grid step to $\Delta l=1$ nm within this region. To reduce computational costs, the symmetry planes $XZ$ and $YZ$ were utilized for the feeder system, and $YZ$ symmetry was exploited in simulations including the CR. The dipole source emitted a short pulse covering the spectral region between $300$ and $800$ nm. The total duration of the simulations was set to $T=150$ fs. The analyses focused on a target wavelength $\lambda_{Ag}=450$ nm for silver and $\lambda_{Au}=600$ nm for gold, both red-shifted from their respective resonances with similar spectral separations.
For the hybrid nano-antenna in an homogeneous dielectric environment, the enhancement factor of the radiative decay rate, $f_{R}$, due to the Purcell effect [22] can be calculated using a fully classical model [21] as the ratio between the total power collected in the far-field region when the noble metal antenna is present, $P_{rad}$, and the power that would be radiated from a point dipole source in free space, $P_{0}$,
@@
where $\gamma_{r}$ ($\gamma_{r,0}$) is the radiative emission rate in the presence (absence) of the plasmonic structure. $P_{rad}$ and the radiative emission rate enhancement factor are readily obtained from FDTD computations [23, 9]. For this purpose, the atom-like emitter is modeled using a time-windowed dipole soft source. Instead of setting a given value of the electromagnetic field at a point in space, the field from this type of source adds to the existing electromagnetic fields without causing scattering, therefore permitting the modeling of fluorescence. The value of $P_{0}$ can be computed from its analytical expression [24], but it is also provided by Lumerical via the sourcepower script command.
@@
$P_{loss}$ is computed as $P_{loss}=P-P_{rad}$ where $P$ is the power radiated by the dipole in the NA environment used to modify its the radiation in vacuum, $P_{0}$.
The total photon emission rate is influenced by radiative and non-radiative transitions and losses in the noble metal structures. The emission yield is characterized by quantum efficiency $\eta_{Q}$ of the process: the ratio between the rate of radiative transitions and the total transition rate including radiative $\gamma_{r}$ and non-radiative $\gamma_{nr}$ transitions and the rate of photons lost in the metals $\gamma_{l}$ [21],
@@
$\eta_{Q}\simeq\dfrac{\gamma_{r}}{\gamma_{r}+\gamma_{l}}=\dfrac{f_{R}}{f_{R}+f_{L}}$ (4)
where we have divided both numerator and denominator of the above expression by $\gamma_{r,0}$. If we substitute (1) and (2) into (4) the resulting expression coincides with the classical definition of the radiation efficiency of an antenna [15, 16]
$\eta=\dfrac{P_{rad}}{P_{rad}+P_{loss}}$ (5)
The evaluation of the emission properties for the NAs in this study is based on the calculation of $f_{R}$ and $f_{L}$ from the FDTD results to gauge the photon emission rate. Both gold and silver plasmonic NA were considered, with their respective optical constants obtained from [25] and [26].
The external photon collection capabilities of the hybrid system are evaluated using of the on-axis directivity, determined from the far-field radiation pattern as
@@
where $K_{0}=K\left(\theta=\pi/2,\phi=\pi/2\right)$ is the radiation intensity emitted along the $Y$ axis, i.e., the power per unit solid angle in that direction. In Eq. (6), $D$ measures the ratio of the power radiated per unit solid angle along the $Y$-axis to that of a hypothetical isotropic radiator [15, 16].
@@
combines the emission efficiency $\eta$ and directivity $D$, serving as the key metric for antenna performance in a RF link [15, 16]. This figure of merit is applicable to optical NAs, where it indicates the potential for photon collection from single-photon emitters. We now analyze the values of $D$, $f_{R}$ and $f_{L}$, the three parameters that determine $G$ for silver and gold NAs.
@@ 3 Results and Discussion
An initial analysis was performed on the isolated dipole-NP feeder problem. This permits the full evaluation of the impact on the emitter properties of the incorporation of the CR. To this end, a series of FDTD calculations were performed on the geometry displayed on Fig. 1 in the absence of the metal sheets defining the CR.
The resonances in the spectral responses of NPs are responsible for their characteristic colors and their intensities and frequencies depend on the particle size, shape and the dielectric properties of the surrounding medium. Fig. 2 displays the resonances calculated from the emission spectrum of the corresponding NPs coupled to the dipole emitter. For gold NPs surrounded by air, the calculations show a negligible displacement of the plasmon resonance for the range of sizes considered in accordance with previous calculations [27]. Furthermore, The spectral shift of the resonance with the particle size correlate well with the red-shift of the resonance as the particle size is increased observed in measurements performed for colloidal suspensions of silver NPs of sizes varying between $60$ and $80$ nm in [28, 29].
@@
Figure 2: Calculated resonance frequencies of silver (a) and gold (b) NPs as a function of their diameter.
Fig. 3 displays the results obtained for the normalized factors $f_{R}$ and $f_{L}$ obtained from FDTD calculations for silver NPs of diameter $d$ coupled to a dipole emitter at a distance $s$ and perpendicular to the spherical NP. Particle sizes between $d=30$ and $80$ nm have been considered and dipole separations ranging from $s=5$ nm to $s=30$ nm. The plots display data collected at the resonance and at the target wavelength of $\lambda_{Ag}=450$ nm. The corresponding results for gold NPs with target wavelength at $\lambda_{Ag}=600$ nm are shown in Fig. 4. In this case, the calculations have been performed for the same range of parameters $d$ and $s$.
@@
Figure 3: Radiative transition rate enhancement, $f_{R}$, and loss, $f_{L}$, factors of silver NPs calculated for different spherical particle diameter $d$ and separation $s$ from the the dipole emitter. $(a)$ and $(c)$ plots are the values at resonance and $b$ and $d$ at the target wavelength of $\lambda_{Ag}=450$ nm.
@@
Figure 4: Radiative transition rate enhancement, $f_{R}$, and loss, $f_{L}$, factors of gold NPs calculated for different spherical particle diameter $d$ and separation $s$ from the the dipole emitter. $(a)$ and $(c)$ plots are the values at resonance and $b$ and $d$ at the target wavelength of $\lambda_{Au}=600$ nm.
@@
The calculated values of the radiative transition factor, $f_{R}$, at the target wavelengths generally increase as the particle size grows and the distance between the emitter and the NP diminishes, both for silver and gold NPs. However, the variation of $f_{R}$ with particle size $d$ at resonance differs between silver and gold. For gold NPs, the trend in $f_{R}$ at resonance is similar to that observed at $\lambda_{Au}$, while for silver NPs, $f_{R}$ exhibits a local maximum at a specific $d$ when the separation $s$ is small. In general, $f_{R}$ reaches higher values for silver NPs than for gold, both at resonance and at the target wavelength. Although the relative size of the NPs compared to the wavelength at which $f_{R}$ is measured is larger for silver NPs, the ratio of the relative sizes is smaller than the observed ratios of $f_{R}$ values. On the other hand, when the value of the wavelength is taken in consideration, the corresponding relative distances between the source and the NP are larger for silver NPs.
@@
Regarding photon loss due to absorption at the NP, the value of $f_{L}$ increases significantly as the distance between the dipole emitter and the NP decreases. Conversely, the NP size $d$ shows a lesser effect on $f_{L}$. Overall, the loss factors for silver NPs are smaller, though still comparable, to those for gold NPs.
@@
As discussed in Section 2, achieving efficient emission requires a good balance between relatively low $f_{L}$ values and sufficiently high $f_{R}$ values. Based on the analysis of the coupled dipole-nanoparticle system, a particle size of $d=70$ nm with a separation of $s=15$ nm from the dipole was chosen for the analysis of the full structure incorporating the CR. For these parameters, at their respective target wavelengths, the results show $f_{L}=5.4$ and $f_{R}=8.3$ for silver nanoparticles, and $f_{L}=6.1$ and $f_{R}=4.9$ for gold nanoparticles. This indicates a higher radiative emission rate enhancement with lower losses for silver compared to gold.
@@
Under these conditions, $f_{R}>f_{L}$ for silver NPs whereas $f_{R}<f_{L}$ for gold NPs. The corresponding values of the emission gain figure of merit, $G$, as defined in (7), with a directivity value of $D=1.5$ for an electrically short dipole [15, 16], is $G_{Ag}=0.9$ for silver and $G_{Au}=0.7$ for gold NPs.
@@
The radiation properties of CRs in RF antennas were systematically studied in [30] through extensive experimental measurements of the gain achieved with different corner configurations using a resonant half-wave dipole feeder. As the distance from the corner apex to the feeder increases, a series of positions corresponding to local maxima of the antenna gain are identified. The first three positions were thoroughly analyzed in [30].
@@
Whereas the second and third positions are sharply defined, the first position corresponds to a broader region. This is due to a trade-off between the increase in directivity obtained by moving the feeder closer to the apex and the associated rise in ohmic losses in the conductors. In the first position, the corner angle for maximum gain depends on the width and length of the reflecting planes: as the planes become larger, the optimal angle decreases and the maximum gain increases. Conversely, fixed optimal values of the corner angle are observed for the second and third positions when the planes are sufficiently large. For the second position, a maximum gain occurs at $\psi=65^{\circ}$ when the width and length of the corner planes exceed about one wavelength. At the third position, an optimal corner angle is found at $\psi=90^{\circ}$ when the reflecting planes are larger than half a wavelength.
@@
In the following analyses, we select a corner angle of $\psi=65^{\circ}$ and corner plane dimensions of $L=1.5$ $\mu$m and $H=1$ $\mu$m. We examine the directivity $D$ along the corner axis and investigate how the presence of the CR influences the values of $f_{R}$ and $f_{L}$ at the target wavelengths for silver and gold NAs.
@@
Fig. 5 presents the on-axis directivities calculated for silver (a) and gold (b) CRs, with film thicknesses ranging from $W=10$ nm to $80$ nm and feeder positions between $S=0.2$ $\mu$m and $S=1$ $\mu$m. For the presentation of the results, the values of $S$ have been rescaled with the respective target wavelengths for the gold and silver reflectors at which the directivities are evaluated.
The figure reveals similar patterns in the variation of directivity with the normalized separation $S/\lambda_{M}$ and the film thickness $W$, where $M$ denotes the corresponding noble metal. The higher peak directivities obtained for silver than for gold can be attributed mainly to the larger relative values of the corner plane length $L/\lambda_{Ag}>L/\lambda_{Au}$.
The metal film thickness also plays a significant role in the radiation properties of the NA. As the thickness increases, the peak directivity increases, and the variation with $S/\lambda_{M}$ follows a more regular oscillatory pattern with local maxima and minima as in the case of RF antennas [30]. The irregularities in variations of directivity at smaller $W$ are less pronounced for gold mirrors compared to silver. Additionally, the positions of maxima and minima shift to larger values of $S/\lambda_{M}$ as $W$ increases. At the largest thickness of $W=80$ nm, there is directivity null precisely at $S=\lambda_{M}$.
A better understanding of the variations observed in the curves of $D$ can be achieved through a detailed analysis of the radiation patterns in each case. Such plots are presented for silver CRs with $W=10$ nm, $W=40$ nm and $W=80$ nm ,respectively, in figures 6, 7 and 8.

### v2

# Corner Reflector Plasmonic Nanoantennas for Enhanced Single-Photon Emission

@@ Abstract
The emission rate of atom-like photon sources can be significantly improved by coupling them to plasmonic resonant nanostructures. These arrangements function as nanoantennas, serving the dual purpose of enhancing light–matter interactions and decoupling the emitted photons. However, there is a contradiction between the requirements for these two tasks. A small resonator volume is necessary for maximizing interaction efficiency, while a large antenna mode volume is essential to achieve high emission directivity. In this work, we analyze a hybrid structure composed of a noble metal plasmonic resonant nanoparticle coupled to the atom-like emitter, which is designed to enhance the emission rate, alongside a corner reflector aimed at optimizing the angular distribution of the emitted photons. A comprehensive numerical study of silver and gold corner reflector nanoantennas, employing the finite difference time domain method, is presented. The results demonstrate that a well-designed corner reflector can significantly enhance photon emission directivity while also substantially boosting the emission rate.
@@ 1 Introduction
Efficient and reliable single-photon sources, capable of producing on-demand high-purity single-photon states, are key elements within the emerging field of photon-based quantum technologies [1, 2]. Atom-like emitters, based either on solid-state or molecular systems, are a promising alternative for this purpose [3, 4, 5, 6, 7]. In the past years, nanoscale elements have been introduced to enhance the emission rates of photon sources and to tailor their spatial radiation properties. Such nanoantennas (NAs) reproduce at the nanoscale the action of traditional radio frequency (RF) aerials. Various traditional antenna geometries have been considered for applications in the optical realm: dipole [8, 9], patch [10], bowtie [11] or the Yagui–Uda array [12, 13, 14].
@@
Antennas for light [12] serve the dual purpose of enhancing emitter fluorescence and facilitating the outcoupling of the generated photons. The performance requisites for the two functions are antithetical: whereas a small mode volume is required for an enhanced emission rate, high antenna directivity is well known to be the result of a large electrical size, i.e, the extent of the structure measured using the wavelength of the electromagnetic field as unit reference [15, 16]. In [4], such antipodal requirements have been reconciled by photomodifying a nanopatch antenna assembled around a nanodiamond with a nitrogen-vacancy center to reduce the cavity mode size while keeping a large size emitter to spatially shape the radiation produced. In this work, we tackle this issue by using a hybrid plasmonic structure built from the combination of a conventional NA acting as the feeder element of a corner reflector (CR) [17]. This strategy simultaneously permits optimization of the emission rate at the feeder design and the capacity to use a CR to enhance the directivity of the photons produced.
@@
The use of CRs in antenna design has been prevalent since Kraus introduced them in 1940 [17], owing to their outstanding performance, simplicity, and ease of construction. This work demonstrates that these advantageous properties can also be applied to nanoscale photonics. The potential of CRs to enhance photon emission was highlighted in [18], where the radiation patterns generated by an electric dipole emitter with a perfect conductor CR were analyzed. Here, we present a thorough numerical study of the proposed NA, utilizing two alternative noble metals: silver and gold. The feeder is assumed to be constituted by an atom-like dipole emitter coupled to a spherical nanoparticle (NP), though the proposed CRs can also accommodate more complex feeder geometries, such as dimers or bowtie structures [8, 11], which may offer enhanced emission. The analysis is conducted using the finite difference time domain method (FDTD) [19, 20], enabling the calculation of the radiative emission rate enhancement factor, losses within the metallic structures, and the far-field patterns of the emitted photons. This approach provides a comprehensive characterization of the NA.
@@
The geometry employed in the analyses is depicted in Figure 1. A spherical NP with diameter $d$ is positioned at the origin. The atom-like emitter is represented by a point dipole source $\mathbf{p}$ located at a distance $s$ from the surface of the metal nanosphere that is placed at a distance $S$ from the corner apex along the bisector. The corner angle is $\psi$, while the metal films have a length $L$ and height $H$, as shown in the figure. The thickness of the metal films forming the reflector is $W$. The planes of the CR are assumed to be made of the same metal as used for the NP in the feeder system.
@@
A metallic plane typically reflects a substantial portion of an incident wave originating from a large distance, due to the significant mismatch between its optical properties and those of the surrounding dielectric. However, when a noble metal plane is placed near a dipolar source, it can significantly influence the emission rate [21] in a similar manner to the behavior observed with emission near NPs. Therefore, analyzing only the radiation pattern of emitted photons is insufficient for designing CR-based NAs. A comprehensive analysis, including the effects on radiative emission and loss rates, has been conducted. The impact of the thickness of the metal films on these properties has also been addressed.
@@ 2 Materials and Methods
The numerical calculations were conducted using ANSYS Lumerical FDTD Solutions, release 2024 R2.1, on a HP Z6 G4 Workstation with two Intel Xeon 6226R processors and 512 GB of RAM. For the FDTD simulations, a variable-size discretization was employed, with a finer grid in the dipole-nanoparticle region. The discretization parameters were adjusted based on a series of convergence analyses, setting the maximum grid step to $\Delta l=1$ nm within this region. To reduce computational costs, the symmetry planes $XZ$ and $YZ$ were utilized for the feeder system, and $YZ$ symmetry was exploited in simulations, including the CR. In principle, exploiting the symmetry of the problem when specifying the boundary conditions should not affect the accuracy of the obtained solutions, which should remain consistent with the symmetry considerations, regardless of the boundary conditions used. Nevertheless, several test computations were carried out without the symmetry boundary conditions, and no significant differences were observed in the results. The dipole source emitted a short pulse covering the spectral region between $300$ and $800$ nm. The total duration of the simulations was set to $T=150$ fs. The analyses focused on a target wavelength $\lambda_{Ag}=450$ nm for silver and $\lambda_{Au}=600$ nm for gold. The wavelength of interest in any application will depend on the specific fluorophore being used. To ensure a fair comparison between the two metals, which have different plasmonic resonance wavelengths, the target wavelengths for the silver and gold NAs were arbitrarily selected to be redshifted from their respective resonances by similar spectral separations, typical of those found in applications.
In spite of the intrinsically quantum nature of the fluorescence effect, the enhancement factor of the radiative decay rate, $f_{R}$, due to the Purcell effect [22] of a nanoantenna in an homogeneous dielectric environment can be calculated using a fully classical model [21, 23] as the ratio between the total power collected in the far-field region when the noble metal antenna is present, $P_{rad}$, and the power that would be radiated from a point dipole source in free space, $P_{0}$:
@@
where $\gamma_{r}$ ($\gamma_{r,0}$) is the radiative emission rate in the presence (absence) of the plasmonic structure. $P_{rad}$ and the radiative emission rate enhancement factor, or Purcell factor, are readily obtained from FDTD computations [24, 9]. For this purpose, the atom-like emitter was modeled using a time-windowed dipole soft source. Instead of setting a given value of the electromagnetic field at a point in space, the field from this type of source adds to the existing electromagnetic fields without causing scattering, therefore permitting the modeling of fluorescence. The value of $P_{0}$ can be computed from its analytical expression [25], but it is also provided by Lumerical via the sourcepower script command. On the other hand, complex non-exponential or non-Markovian emission processes cannot be accounted for by the model.
@@
$P_{loss}$ is computed as $P_{loss}=P-P_{rad}$, where $P$ is the power radiated by the dipole in the NA environment used to modify its radiation in a vacuum—$P_{0}$.
The total photon emission rate is influenced by radiative and non-radiative transitions and losses in the noble metal structures. The emission yield is characterized by the quantum efficiency $\eta_{Q}$ of the process: the ratio between the rate of radiative transitions and the total transition rate, including radiative $\gamma_{r}$ and non-radiative $\gamma_{nr}$ transitions, and the rate of photons lost in the metals $\gamma_{l}$ [21],
@@
$\eta_{Q}\simeq\dfrac{\gamma_{r}}{\gamma_{r}+\gamma_{l}}=\dfrac{f_{R}}{f_{R}+f_{L}},$ (4)
where we have divided both numerator and denominator of the above expression by $\gamma_{r,0}$. If we substitute (1) and (2) into (4), the resulting expression coincides with the classical definition of the radiation efficiency of an antenna [15, 16]:
$\eta=\dfrac{P_{rad}}{P_{rad}+P_{loss}}.$ (5)
The evaluation of the emission properties for the NAs in this study is based on the calculation of $f_{R}$ and $f_{L}$ from the FDTD results to gauge the photon emission rate. Both gold and silver plasmonic NA were considered, with their respective optical constants obtained from [26, 27].
The external photon collection capabilities of the hybrid system were evaluated using on-axis directivity, which is determined from the far-field radiation pattern as
@@
where $K_{0}=K\left(\theta=\pi/2,\phi=\pi/2\right)$ is the radiation intensity emitted along the $Y$ axis, i.e., the power-per-unit solid angle in that direction. In Equation (6), $D$ measures the ratio of the power radiated per unit solid angle along the $Y$ axis to that of a hypothetical isotropic radiator [15, 16].
@@
combines the emission efficiency $\eta$ and directivity $D$, serving as the key metric for antenna performance in an RF link [15, 16]. This figure of merit is applicable to optical NAs, where it indicates the potential for photon collection from single-photon emitters. We now analyze the values of $D$, $f_{R}$, and $f_{L}$, which are the three parameters that determine $G$ for silver and gold NAs.
The Kildal limit [28] sets a theoretical upper bound on the maximum directivity that can be achieved with an antenna
$D_{max}=(kR)^{2}+3,$ (8)
where $k$ is the wavenumber, and $R$ is the radius of the sphere circumscribing the antenna. Antennas having directivities beyond the limits imposed by their physical dimensions are often named superdirective antennas [29]. It was early realized that the implementation of such antennas is subject to physical restrictions [30, 31] such us very high values of $Q$, i.e., very low bandwidths, that are also associated with very low efficiencies due to very large ohmic losses, as well as very sensitive implementation tolerances [29].
Antenna designs based on dielectric resonators coupled to a dipole emitter have proven particularly advantageous in the pursuit of electrically small, practical superdirective antennas with high efficiency. This design has been theoretically analyzed at the nanoscale [32] and experimentally demonstrated in the UHF band [33]. It is also well established that superdirectivity is not viable for small plasmonic nanoparticles [32]; however, in principle, plasmonic nanostructures can outperform their dielectric counterparts despite the intrinsic metal losses they exhibit [34]. The approach in this work aligns with classical antenna design, where using a reflector provides a straightforward solution to spatially shaping the emitted field into a highly directive beam by enlarging the antenna dimension $R$.
@@ 3 Results and Discussion
An initial analysis was performed on the isolated dipole-NP feeder problem. This permitted the full evaluation of the impact on the emitter properties of the incorporation of the CR. To this end, a series of FDTD calculations were performed on the geometry displayed in Figure 1 in the absence of the metal sheets defining the CR.
The resonances in the spectral responses of NPs are responsible for their characteristic colors, and their intensities and frequencies depend on the particle size, shape, and the dielectric properties of the surrounding medium. Figure 2 displays the resonances calculated from the emission spectrum of the corresponding NPs coupled to the dipole emitter. For gold NPs surrounded by air, the calculations show a negligible displacement of the plasmon resonance for the range of sizes considered, which is in accordance with previous calculations [35]. Furthermore, the spectral shift of the resonance with the particle size correlates well with the redshift of the resonance, as increased particle size was observed in measurements performed for colloidal suspensions of silver NPs of sizes varying between $60$ and $80$ nm in [36, 37]. To obtain the data in Figure 2, the smooth spectra from FDTD simulations were Nyquist interpolated to increase the spectral resolution without altering the information contained in the original data.
@@
Figure 2: Calculated resonance wavelengths of silver (a) and gold (b) NPs as a function of their diameters.
Figure 3 displays the results obtained for the normalized factors $f_{R}$ and $f_{L}$ obtained from FDTD calculations for silver NPs of diameter $d$ coupled to a dipole emitter at a distance $s$ and perpendicular to a spherical NP. Particle sizes between $d=30$ and $80$ nm were considered, as well as dipole separations ranging from $s=5$ nm to $s=30$ nm. The plots display data collected at the resonance and at the target wavelength of $\lambda_{Ag}=450$ nm. The corresponding results for gold NPs with their target wavelength at $\lambda_{Au}=600$ nm are shown in Figure 4. In this case, the calculations were performed for the same range of parameters $d$ and $s$.
@@
Figure 3: Purcell, $f_{R}$, and loss, $f_{L}$, factors of silver NPs calculated for different spherical particle diameter values $d$ and separation values $s$ from the the dipole emitter. (a,c) plots are the values at resonance and, (b,d) are at the target wavelength of $\lambda_{Ag}=450$ nm.
@@
Figure 4: Purcell, $f_{R}$, and loss, $f_{L}$, factors of gold NPs calculated for different spherical particle diameter values $d$ and separation values $s$ from the the dipole emitter. (a,c) plots are the values at resonance, and (b,d) are at the target wavelength of $\lambda_{Au}=600$ nm.
@@
The calculated values of the radiative transition factor, $f_{R}$, at the target wavelengths generally increased as the particle size grew and the distance between the emitter and the NP diminished, both for silver and gold NPs. However, the variation in terms of $f_{R}$ with particle size $d$ at resonance differed between silver and gold. For gold NPs, the trend for $f_{R}$ at resonance was similar to that observed at $\lambda_{Au}$, while for silver NPs, $f_{R}$ exhibited a local maximum at a specific $d$ when the separation $s$ was small. In general, $f_{R}$ reached higher values for silver NPs than for gold, both at resonance and at the target wavelength. Although the relative size of the NPs compared to the wavelength at which $f_{R}$ was measured was larger for silver NPs, the ratio of the relative sizes was smaller than the observed ratios of $f_{R}$ values. On the other hand, when the value of the wavelength was taken in consideration, the corresponding relative distances between the source and the NP were larger for silver NPs.
@@
Regarding photon loss due to absorption at the NP, the value of $f_{L}$ increased significantly as the distance between the dipole emitter and the NP decreased. Conversely, the NP size $d$ showed a lesser effect on $f_{L}$. Overall, the loss factors for silver NPs were smaller, though still comparable, to those for gold NPs.
@@
As discussed in Section 2, achieving efficient emission requires a good balance between relatively low $f_{L}$ values and sufficiently high $f_{R}$ values. Based on the analysis of the coupled dipole-nanoparticle system, a particle size of $d=70$ nm with a separation of $s=15$ nm from the dipole was chosen for the analysis of the full structure incorporating the CR. For these parameters, at their respective target wavelengths, the results show $f_{L}=5.4$ and $f_{R}=8.3$ for silver nanoparticles and $f_{L}=6.1$ and $f_{R}=4.9$ for gold nanoparticles. This indicates a higher radiative emission rate enhancement with lower losses for silver compared to gold.
@@
Under these conditions, $f_{R}>f_{L}$ for silver NPs, whereas $f_{R}<f_{L}$ for gold NPs. The corresponding values of the emission gain figure of merit, $G$, as defined in (7)—with a directivity value of $D=1.5$ for an electrically short dipole [15, 16]—were $G_{Ag}=0.9$ for silver and $G_{Au}=0.7$ for gold NPs.
@@
The radiation properties of the CRs in RF antennas were systematically studied in [38] through extensive experimental measurements of the gain achieved with different corner configurations using a resonant half-wave dipole feeder. As the distance from the corner apex to the feeder increased, a series of positions corresponding to local maxima of the antenna gain was identified. The first three positions were thoroughly analyzed in [38].
@@
Whereas the second and third positions were sharply defined, the first position corresponded to a broader region. This is due to a trade-off between the increase in directivity obtained by moving the feeder closer to the apex and the associated rise in ohmic losses in the conductors. In the first position, the corner angle for maximum gain depended on the width and length of the reflecting planes: as the planes became larger, the optimal angle decreased and the maximum gain increased. Conversely, fixed optimal values of the corner angle were observed for the second and third positions when the planes were sufficiently large. For the second position, a maximum gain occurred at $\psi=65^{\circ}$ when the width and length of the corner planes exceeded about one wavelength. At the third position, an optimal corner angle was found at $\psi=90^{\circ}$ when the reflecting planes were larger than half a wavelength.
@@
In the following analyses, we selected a corner angle of $\psi=65^{\circ}$ and corner plane dimensions of $L=1.5$ $\upmu$m and $H=1$ $\upmu$m. We examined the directivity $D$ along the corner axis and investigated how the presence of the CR influenced the values of $f_{R}$ and $f_{L}$ at the target wavelengths for silver and gold NAs.
Figure 5 presents the on-axis directivities calculated for silver (a) and gold (b) CRs, with film thicknesses ranging from $W=10$ nm to $80$ nm and feeder positions between $S=0.2$ $\upmu$m and $S=1$ $\upmu$m. For the presentation of the results, the values of $S$ have been rescaled with the respective target wavelengths for the gold and silver reflectors at which the directivities were evaluated.
@@
The figure reveals similar patterns in the variation in directivity with the normalized separation $S/\lambda_{M}$ and the film thickness $W$, where $M$ denotes the corresponding noble metal. The higher peak directivities obtained for silver than for gold can be attributed mainly to the larger relative values of the corner plane length $L/\lambda_{Ag}>L/\lambda_{Au}$.
The metal film thickness also played a significant role in the radiation properties of the NA. As the thickness increased, the peak directivity increased, and the variation in $S/\lambda_{M}$ followed a more regular oscillatory pattern in terms of the local maxima and minima, as in the case of RF antennas [38]. The irregularities in variations of directivity at smaller $W$ values were less pronounced for gold mirrors compared to silver. Additionally, the positions of the maxima and minima shifted to larger values of $S/\lambda_{M}$ as $W$ increased. At the largest thickness of $W=80$ nm, the directivity was null precisely at $S=\lambda_{M}$.
A better understanding of the variations observed in the curves of $D$ can be achieved through a detailed analysis of the radiation patterns in each case. Such plots are presented for silver CRs with $W=10$ nm, $W=40$ nm, and $W=80$ nm, respectively, in Figures 68.
One of the main differences observed at small film thickness ($W=10$ nm) was the presence of relevant amplitude side lobes at the lower values of $S$. This phenomenon led to reduced directivity values, despite the main lobe exhibiting an narrower beamwidth at $W=10$ nm. Conversely, the patterns seen at $W=40$ nm and $W=80$ nm were qualitatively similar, indicating a convergence as the thickness $W$ increased.
For thicker-sheet CRs, the gradual narrowing (or broadening) of the on-axis lobe contributed to regions of increasing (or decreasing) directivity values. The zero observed in the directivity curves arose from the splitting of the main lobe at $S=\lambda_{M}$ that produced a radiation null along the $Y$ axis. In contrast, for $W=10$, the local minima in the directivity plots were correlated with the increasing relative weight of the side lobes in the radiation pattern. As $S$ increased, separation of the main lobe was also noted; however, it occurred along the horizontal plane, unlike the vertical splitting seen at $W=40$ nm and $W=80$ nm. As $S$ continued to grow, additional lobe bifurcations became apparent, leading to the emergence of a fractal structure in the radiation pattern for very large values of $S$ in extensive CRs, as was analyzed in [18].
The sequences of radiation diagrams presented in Figures 68 ended up being very similar for the corresponding ones of gold CRs when the values of $S$ were normalized by the respective operation wavelength $\lambda_{M}$.
We now examine the impact of the corner reflector on the losses and the emission rate of the dipole-NP system. Figure 9 displays the $f_{R}$ and the normalized loss factor $f_{L}$ of NAs incorporating silver CRs of four different thicknesses, $W$, as a function of the normalized distance, $S/\lambda_{Ag}$, between the silver NP and the corner apex. The corresponding results for the gold NAs are shown in Figure 10. In all the figures, the horizontal lines pinpoint the reference values for the NPs without the CRs.